OpenStudy (anonymous):

Y varies directly as the square of x. When x = 4, y=8 . Find x when y = 32

OpenStudy (anonymous):

Is that exactly how the question is written?

OpenStudy (anonymous):

If y varied as the square of x, y would equal x^2, and 4^2 does not equal 8.

OpenStudy (anonymous):

thats what im saying and thats how it is written

zepdrix (zepdrix):

y varies directly as the square of x y=k(x^2) When x=4, y=8 8=k(4^2) Solve for k! :D Then use your new found k value, and plug in the new y coordinate, and solve for x! :)

OpenStudy (anonymous):

@geoffb 'varies directly' means it's a constant multiple of what it's varying with, so you have to multiply by k.

OpenStudy (anonymous):

Thank Cliff. :) I caught that as soon as I read @zepdrix's response. Definitely didn't think it through.